The radius of the Earth increases by 1 m, how will the length of the Earth’s equator increase?
1
Expert's answer
2010-06-09T10:38:56-0400
The earth’s equator is a circle. And the radius of the Earth is the radius of this circle. The length of a circle is expressed by the formula: L = 2*π*R. So if the radius changes by 1 m (it was R and it became (R+1)m), the length of the circle is: L’ = 2*π*(R+1m) = 2*π*R + 2*π*1m = L + (2*π)m. So the length of the Earth’s equator increases by (2*π)m.